# LLMR: Real-time Prompting of Interactive Worlds using Large Language Models

### 19 Dec 2023 v2

@@ 8.1.1. Experiment Setup
Figure 13. An illustration of our experimental setup. We provide a view of the bathroom scene (left) and a subset of the 150 prompts (right) used in this space for the evaluation provided in Figure 14.
@@
Figure 14. Comparison for average compilation and run-time error rate. SA stands for the Scene Analyzer, and SL stands for the Skill Library. Overall, in both creating from scratch, as well as editing existing scenes, LLMR outperforms GPT-4 by 3x in the case with few-shot prompting, and gives over 4x improvement compared to zero shot performance of GPT-4.
@@ 8.2.1. Methodology
Figure 15. Performance Improvement of LLMR Modules Organized by Difficulty Level of Prompts Comparing Error Rates of GPT-4 with Incremental LLMR Module Integration
@@ 8.3. Task Complexity
Specific Skill Requirement – Certain tasks are inherently more difficult. For example, deforming the mesh of an object is much more complicated than adding an object to the scene. A human developer may need to look up examples and documentation to achieve a complex task; LLMR can reduce the complexity of the task by starting a template. However, LLMR is not error-proof. As shown in the previous section, LLMR’s error rate increases (but not above 40%) as a task becomes more difficult. The error rate can be further brought down by adding relevant skills to the Skill Library by an experienced developer to help LLMR achieve a higher success rate and reduce possible rounds of iteration between the Builder and Inspector. LLMR can save the time of experienced developers by generalizing beyond the examples provided in the Skill Library.
Token Requirement – The amount of tokens required grows as the scene or the object becomes more complex. For example, if the existing scene has a tree object with many leaves, where each leaf is considered a child game object, the scene summary could easily exceed the maximum token allowed (at the time of writing, the maximum number of tokens was 8k, though this has now grown significantly). In anticipation of this, our Scene Analyzer module only fetches the top-level game object name to filter out the relevant game object based on user query and task. This allows our framework to handle a scene as complex as the Kitchen scene (example in Figure 10) and the Bathroom scene (used in Numerical Study, Figure 13).
Besides the complexity of objects and scenes, a task itself can also require a lot of tokens. One such example is generating an animation (with the help of a skill written for the Skill Library) that involves generating time-series of numerous joint positions and rotations. An optimization could involve developing a simple way to represent the time-series information.
@@
Quality Requirement – A user may request different levels of fidelity of the output. For example, the user could create a complex scene out of primitives only with the help of the Planner module (e.g., a full kitchen, Appendix 18) instead of out of higher fidelity 3D models (see examples of participants’ creations in the video figure). The flexibility to create visually simple yet functional and interactive scenes is akin to creating a lo-fi mockup that allows users to quickly prototype and iterate without waiting for the full generation of 3D scenes that are visually complex but cost a lot of compute and time and are not easy to modify.
@@ 8.5.1. Methodology
Model
Single prompts, empty [sec]
Single prompts, bathroom [sec]
Sequential prompts [sec]
GPT-4
35.240
20.600
77.1
GPT-4, few-shot
37.820
21.280
112.5
GPT-4, few-shot with SA
33.900
20.580
69.4
GPT-4, few-shot with SA, SL
34.460
21.640
74.6
LLMR
90.980
49.160
170.9
@@ 9.2.1. Approach to Prompting and Instruction Strategies
9.2.2. Comparison with Prior Approach 3D World Creation
@@ 9.2.3. Considerations and Challenges
Participants noted that our framework’s output was often unpredictable compared to traditional methods, and there was uncertainty about whether the model would understand the desired structure (P0). Some participants pointed out that the choice between traditional methods and our generative method of creating 3D worlds might depend on the artistic nature of the project and the need for creative input (P5, P7, P11). Other participants recognized that for projects requiring precise, structured, or rigid control, traditional tools might be preferred over our framework (P4, P8). Lastly, they also mentioned that for more complex tasks or as projects grew in size, manual code editing might still be necessary, as it could be faster than creating detailed descriptions for the model (P1, P4, P8).
@@ 10. Ethical Considerations
While LLMR and other LLM tools can be transformative to many industries and applications, there exist risks with any AI-enabled systems. Firstly, the concern of developers and creators being replaced has been on the surface of discussion. However, these tools have not been proven to achieve end-to-end development. Participants of our user study commented that our framework is better at integrating human intervention and involvement, and thus our framework helps improve productivity and facilitate brainstorming, rather than completely automating the creation process. A more serious concern is the potential for individuals to generate harmful and inappropriate content with our framework. Despite the safeguards put in by Sketchfab and OpenAI through content moderation and model alignment, it is still possible to creatively circumvent these safeguards (Liu et al., 2023). While the Roslyn compiler can automatically check for unsafe code, the need for research on how to moderate 3D content is merited.
@@ 11. Limitations and Future Work
Limitation in scene understanding
Currently, our framework requires access to a ”scene graph” with descriptions and the hierarchy of the game objects. The scene graph provides the spatial relationship of game objects and it assumes that the names of the game objects (and their children game object components) are correct and unique. However, 3D models from repositories like Sketchfab often have random, non-descriptive component names. At the moment, the framework manipulates objects by finding the game objects with the exact name, which is not always reliable.
@@
Incorporating world feedback and direct user feedback
Similar to how the Builder-Inspector loop reduces code compilation error, the framework’s understanding of the world could be further improved by incorporating feedback from the virtual world and from the user. For example, if a 3D model is loaded from Sketchfab, the framework is ignorant of the model’s (and their subcomponents’) orientation and center of pivot, and thus does not consistently produce the desired output when asked to rotate the 3D model.
Limitation in Memory and Traceability
Another limitation of the framework is the token size. As mentioned in section 5, we have optimized access to historical conversation and generation to reduce token usage. There is an inherent tradeoff, where the user instruction might refer to something in a previous prompt exchange that is not exposed to the next exchange. For example, the Scene Analyzer has access to the name of the script, the summary of the script, and the public fields of the script, but if the user just wants to change a specific part of a previous script generated two prompts prior, the framework would not know what to do.
@@
Automatic Skill Generation
At the moment, skills in the Skill Library are created by human users. For example, we incorporated the skill of loading assets from Sketchfab as well as the skill of making objects ”grabbable” using MRTK’s (Microsoft, 2017) namespaces. The ability to automatically generate new skills (Wang et al., 2023d) based on a couple of examples would allow our framework to achieve more complex tasks (such as generating animations) and to be compatible with different platforms (such as Quest and ARKit).
Interoptability
We built upon the Unity engine for its robustness and the large amount of existing examples of C# code that our LLMs have likely seen during training. Our work is independent of Unity and its closed-beta AI tools  [^1] https://unity.com/ai, although our tool can be an add-on to Unity. We want to highlight that our framework can be adapted to any environment that supports run-time compilation. Unity is the baseline requirement for using our framework, and a web-based approach would further make prompt-based interactive 3D worlds easy to share and collaborate within. In fact, some of our user study participants work on web-based mixed reality development, and they commented that our framework can be easily adapted to their coding environment.
@@ Acknowledgements.
We would like to thank various collaborators who gave very useful comments and suggestions, in particular Jennifer Marsman, Jason Carter, Sebastien Vandenberghe, Haiyan Zhang, Nebojsa Jojic, Andy Wilson, Gavin Jancke, Grace Hwang, Lili Cheng, Octavio Martinez, Pavan Davuluri, Robin Seiler, Ruben Caballero, Anuj Gosalia, David Wolf, Marc Pollefeys, Miguel Susffalich, Allan Naim, Darren Bennett, Doug Berrett, Gilles Zunino, Jay Watts, Tucker Burns, Preet Mangat. Additionally, we would like to thank all the study participants.

### 22 Mar 2024 v3

@@ 8.1.2. Results and Discussion
Figure 14. Comparison for average compilation and run-time error rate. SA stands for the Scene Analyzer, SL stands for the Skill Library, and I stands for the Inspector. Overall, in both creating from scratch, as well as editing existing scenes, LLMR outperforms GPT-4 by 3x in the case with few-shot prompting, and gives over 4x improvement compared to the performance of zero-shot GPT-4.
Comparison for average compilation and run-time error rate. The image presents two bar charts side by side, comparing the average compilation and runtime error rates across different methods. Description of the first bar chart, labeled ”Single Prompt, Empty Scene”: The chart displays five bars, each representing different combinations of techniques and prompting methods with GPT-4. The bars are shades of grey and orange, with the orange bars representing the LLMR (GPT-4, few-shot, SA, SL) method. Each bar is annotated with a mean (M) value at the top and a standard error (SE) value just below it. Description of the second bar chart, labeled ”Single Prompt, Existing Scene”: Similar to the first chart, it has six bars with the same color scheme and annotations.
Figure 14. Comparison for average compilation and run-time error rate. SA stands for the Scene Analyzer, SL stands for the Skill Library, and I stands for the Inspector. Overall, in both creating from scratch, as well as editing existing scenes, LLMR outperforms GPT-4 by 3x in the case with few-shot prompting, and gives over 4x improvement compared to the performance of zero-shot GPT-4.
@@ 8.2.2. Results and Discussion
Figure 15. Performance Improvement of LLMR Modules Organized by Difficulty Level of Prompts. Comparing Error Rates of GPT-4 with Incremental LLMR Module Integration. The error rates for most methods increase with difficulty, and the LLMR method (in orange) still maintains a consistently lower error rate compared to others.
Comparison for average compilation and run-time error rate. The image contains two line graphs placed side by side. Description of the first graph, labeled ”Empty Scene”: This graph plots the average error rate on the y-axis, which ranges from 0 to 0.7, against varying levels of difficulty on the x-axis, labeled from ”Easy (1-2)” to ”Hard(10)”. There are four lines representing different methods: GPT-4 Zero Shot, GPT-4 + Few Shot, GPT-4 + Few Shot + Scene Analyzer, and LLMR (GPT-4 + Few Shot + Scene Analyzer + Inspector + Skill Library), each marked in different colors and with corresponding symbols. The LLMR line (in orange) demonstrates a generally lower error rate across all difficulty levels. Description of the second graph, labeled ”Existing Bathroom Scene”: This graph has a similar structure, with the average error rate on the y-axis and difficulty on the x-axis. It shows the performance of the same four methods. In this scenario, the error rates for all methods increase with difficulty, but the LLMR method (in orange) still maintains a lower error rate compared to others, although the differences are less pronounced compared to the ”Empty Scene” graph.
Figure 15. Performance Improvement of LLMR Modules Organized by Difficulty Level of Prompts. Comparing Error Rates of GPT-4 with Incremental LLMR Module Integration. The error rates for most methods increase with difficulty, and the LLMR method (in orange) still maintains a consistently lower error rate compared to others.
@@ 8.3. Task Complexity
Specific Skill Requirement – Certain tasks are inherently more difficult. For example, deforming the mesh of an object is much more complicated than adding an object to the scene. A human developer may need to look up examples and documentation to achieve a complex task; LLMR can reduce the complexity of the task by starting a templated script. However, LLMR is not error-proof. As shown in the previous section, LLMR’s error rate increases (but not above 40%) as a task becomes more difficult. The error rate can be further brought down by adding relevant skills to the Skill Library by an experienced developer to help LLMR achieve a higher success rate and reduce possible rounds of iteration between the Builder and Inspector. LLMR can save the time of experienced developers by generalizing beyond the examples provided in the Skill Library.
Token Requirement – The amount of tokens required grows as the scene or the object becomes more complex. For example, if the existing scene has a tree object with many leaves, where each leaf is considered a child game object, the scene summary could easily exceed the maximum token allowed (at the time of writing, the maximum number of tokens was 8k, though this has now grown significantly). In anticipation of this, our Scene Analyzer module only fetches the top-level game object name to filter out the relevant game object based on user query and task. This allows our framework to handle a scene as complex as the Kitchen scene (example in Figure 10) and the Bathroom scene (used in Numerical Study, Figure 13). Besides the complexity of objects and scenes, a task itself can also require a lot of tokens. One such example is generating an animation (Huang et al., 2023a) (with the help of a skill written for the Skill Library) that involves generating time-series of numerous joint positions and rotations. An optimization could involve developing a simple way to represent the time-series information.
@@
Quality Requirement – A user may request different levels of fidelity of the output. For example, the user could create a complex scene out of primitives only with the help of the Planner module (e.g., a full kitchen, Figure 1) instead of out of higher fidelity 3D models (see examples of participants’ creations in the video figure). The flexibility to create visually simple yet functional and interactive scenes is akin to creating a lo-fi mockup that allows users to quickly prototype and iterate without waiting for the full generation of 3D scenes that are visually complex but cost a lot of compute and time and are not easy to modify.
Model
@@
Model
Single prompts, empty [sec]
Single prompts, bathroom [sec]
Sequential prompts [sec]
GPT-4
35.24
20.60
77.10
GPT-4, few-shot
37.82
21.28
112.50
GPT-4, few-shot with SA
33.90
20.58
69.40
GPT-4, few-shot with SA, SL
34.46
21.64
74.60
LLMR
90.98 49.16 170.90
@@ 8.5.2. Results and Discussion
Results from user study. The image depicts a bar chart evaluating user satisfaction with Unity software, differentiated between beginner and experienced users across various criteria. Description: The chart has a 1 to 7 scale, where 1 stands for ’Strongly Disagree’ and 7 for ’Strongly Agree’. There are seven pairs of bars, each representing a different aspect of user experience such as ’Frustrated’, ’Satisfied’, ’Iterative’, ’Achieved Goal’, ’Intuitive’, ’Will Recommend’, and ’Will Use’. Each aspect is evaluated by two bars, one for beginner users (lighter shade) and one for experienced users (darker shade). The bars range in height, reflecting the mean score given by the users, with error bars indicating the standard error. For example, on ’Satisfied’, beginners scored around 5.5 and experienced users scored slightly lower, near 5.3. The ’Will Recommend’ aspect has one of the highest scores for both groups.
@@ 9.2.1. Approach to Prompting and Instruction Strategies
9.2.2. Comparison with Prior Approach to 3D World Creation
@@ 9.2.3. Considerations and Challenges
Participants noted that our framework’s output was often unpredictable compared to traditional methods, and there was uncertainty about whether the model would understand the desired structure (P0). Some participants pointed out that the choice between traditional methods and our generative method of creating 3D worlds might depend on the artistic nature of the project and the need for creative input (P5, P7, P11). Other participants recognized that for projects requiring precise, structured, or rigid control, traditional tools might be preferred over our framework (P4, P8). Lastly, they also mentioned that for more complex tasks or as projects grew in size, manual code editing might still be necessary, as it could be faster than creating detailed descriptions of edits for the model (P1, P4, P8).
@@ 10. Ethical Considerations
While LLMR and other LLM tools can be transformative to many industries and applications, there exist risks with any AI-enabled systems. Firstly, the concern of developers and creators being replaced has been on the surface of discussion. However, these tools have not been proven to achieve end-to-end development. Participants of our user study commented that our framework is better at integrating human intervention and involvement, and thus our framework helps improve productivity and facilitate brainstorming, rather than completely automating the creation process. A more serious concern is the potential for individuals to generate harmful and inappropriate content with our framework. Despite the safeguards put in by Sketchfab and OpenAI through content moderation and model alignment, it is still possible to creatively circumvent these safeguards (Liu et al., 2023a). While the Roslyn compiler can automatically check for unsafe code, the need for research on how to moderate 3D content is merited.
@@ 11. Limitations and Future Work
Limitation in scene understanding – Currently, our framework requires access to a ”scene graph” with descriptions and the hierarchy of the game objects. The scene graph provides the spatial relationship of game objects and it assumes that the names of the game objects (and their children game object components) are correct and unique. However, 3D models from repositories like Sketchfab often have random, non-descriptive component names. At the moment, the framework manipulates objects by finding the game objects with the exact name, which is not always reliable.
@@
Incorporating world feedback and direct user feedback – Similar to how the Builder-Inspector loop reduces code compilation error, the framework’s understanding of the world could be further improved by incorporating feedback from the virtual world and from the user. For example, if a 3D model is loaded from Sketchfab, the framework is ignorant of the model’s (and their subcomponents’) orientation and center of pivot, and thus does not consistently produce the desired output when asked to rotate the 3D model.
Limitation in Memory and Traceability – Another limitation of the framework is the token size. As mentioned in Section 5, we have optimized access to historical conversation and generation to reduce token usage. There is an inherent tradeoff, where the user instruction might refer to something in a previous prompt exchange that is not exposed to the next exchange. For example, the Scene Analyzer has access to the name of the script, the summary of the script, and the public fields of the script, but if the user just wants to change a specific part of a previous script generated two prompts prior, the framework would not know what to do.
@@
Automatic Skill Generation – At the moment, skills in the Skill Library are created by human users. For example, we incorporated the skill of loading assets from Sketchfab as well as the skill of making objects ”grabbable” using MRTK’s (Microsoft, 2017) namespaces. The ability to automatically generate new skills (Wang et al., 2023d) based on a couple of examples would allow our framework to achieve more complex tasks (such as generating animations) and to be compatible with different platforms (such as Quest and ARKit).
Interoptability – We built upon the Unity engine for its robustness and the large amount of existing examples of C# code that our LLMs have likely seen during training. Our work is independent of Unity and its closed-beta AI tools  [^1] Link to Unity’s closed-beta AI tools: https://unity.com/ai, although our tool can be an add-on to Unity. We want to highlight that our framework can be adapted to any environment that supports run-time compilation. Unity is the baseline requirement for using our framework, and a web-based approach would further make prompt-based interactive 3D worlds easy to share and collaborate within. In fact, some of our user study participants work on web-based mixed reality development, and they commented that our framework can be easily adapted to their coding environment.
@@ Acknowledgements.
We would like to thank various collaborators who gave very useful comments and suggestions, in particular Jennifer Marsman, Jason Carter, Sebastien Vandenberghe, Haiyan Zhang, Nebojsa Jojic, Andy Wilson, Gavin Jancke, Grace Huang, Lily Cheng, Octavio Martinez, Pavan Davuluri, Robin Seiler, Ruben Caballero, Anuj Gosalia, David Wolf, Marc Pollefeys, Miguel Susffalich, Allan Naim, Darren Bennett, Doug Berrett, Gilles Zunino, Jay Watts, Tucker Burns, Preet Mangat. Additionally, we would like to thank all the study participants.